# GraspClutter6D: A Large-scale Real-world Dataset for Robust Perception and Grasping in Cluttered Scenes

### v1

@@ Abstract
Robust grasping in cluttered environments remains an open challenge in robotics. While benchmark datasets have significantly advanced deep learning methods, they mainly focus on simplistic scenes with light occlusion and insufficient diversity, limiting their applicability to practical scenarios. We present GraspClutter6D, a large-scale real-world grasping dataset featuring: (1) 1,000 highly cluttered scenes with dense arrangements (14.1 objects/scene, 62.6% occlusion), (2) comprehensive coverage across 200 objects in 75 environment configurations (bins, shelves, and tables) captured using four RGB-D cameras from multiple viewpoints, and (3) rich annotations including 736K 6D object poses and 9.3B feasible robotic grasps for 52K RGB-D images. We benchmark state-of-the-art segmentation, object pose estimation, and grasping detection methods to provide key insights into challenges in cluttered environments. Additionally, we validate the dataset’s effectiveness as a training resource, demonstrating that grasping networks trained on GraspClutter6D significantly outperform those trained on existing datasets in both simulation and real-world experiments. The dataset, toolkit, and annotation tools are publicly available on our project website: https://sites.google.com/view/graspclutter6d.
Index Terms:
Data sets for robotic vision, Data sets for robot learning, Deep learning in grasping and manipulation.I INTRODUCTION
@@
Improved grasping performance: Models trained on our dataset demonstrate significantly better grasping performance than those trained on existing benchmark datasets.
@@ II RELATED WORK
Robotic Grasping Datasets. State-of-the-art robotic grasping methods rely on deep learning approaches, which necessitate large-scale, diverse datasets for robust performance. Early datasets such as Cornell [35] and Jacquard [36] provided valuable resources for grasp learning. However, they were limited to planar grasping in single-object scenes. OCID-Grasp [22] extended this paradigm to cluttered scenes but offered only 75K planar grasp annotations. Synthetic datasets can be scalable alternatives: DexNet [25] with 6.7M planar grasps, ACRONYM [29] with 6-DoF grasps for 8K objects, and MetaGraspNetv2 [14] with 8,000 synthetic scenes. However, they face significant sim-to-real gaps, such as inaccurate sensor noise and contact dynamics.
Among real-world 6-DoF grasp datasets, only a few public datasets exist. GraspNet-1B [7, 18], a widely adopted grasping benchmark, provides 1.1B grasp poses across 190 scenes. Though it established a foundation for 6-DoF grasping research, it remains confined to simple green tabletop environments with relatively low complexity (35.2% occlusion) and no background variation. The recent HouseCat6D dataset [19] enhances diversity with 194 objects, yet remains limited in scale and variation, offering only 10M grasps for 41 table-only scenes with low occlusion (23.5%). GraspClutter6D bridges these gaps by providing 9.3B 6-DoF grasp annotations across 1,000 highly cluttered scenes spanning diverse environments including bins, shelves, and tables.
@@ III-A Data Acquisition
Camera Calibration. We performed extensive calibration to ensure high-quality RGB-D data. We first conducted intrinsic calibration for all cameras using a ChArUco board [40]. Next, relative poses between cameras were obtained using MC-Calib [41], followed by bundle adjustment with PnP-RANSAC [42] applied to matched points [43, 44] on key frames. Camera poses relative to the robot base were determined using ArUco markers [40]. To address systematic depth errors in low-cost RGB-D cameras, we applied depth correction through least-squares fitting with linear models [12]. This reduced mean absolute depth errors from to for the RealSense D415, from to for the RealSense D435, and from to for the Azure Kinect. Our dataset provides undistorted, depth-corrected RGB-D pairs with camera poses, requiring no additional post-processing.
@@ III-B Annotation Pipelines
We first generated integrated point clouds of each scene by merging data from 13 different Zivid camera views. Annotators then aligned 3D object models to these sub-millimeter accurate point clouds, targeting a mean depth error below per object. Our tool visualized depth errors between annotated object models and scene point clouds, allowing annotators to monitor quality and refine alignments through additional adjustment and ICP [45] refinement upon saving each annotation. Independent experts verified all annotations through a double-review process. The verified poses were propagated to all camera views using camera intrinsic parameters and extrinsic poses. This process yielded annotated instances with 6D object poses and segmentation masks. The annotation tool will be made publicly available on our project website.
@@ III-C Dataset Statistics
Dataset Splits. For standardized evaluation, we provide two main dataset configurations: (1) A cross-object setup that tests generalization to novel objects. This split uses 68 unseen YCB-HOPE objects [37, 38] for testing (12K images across 235 scenes), widely used robotic research benchmark objects, and 132 objects for training (32K images across 413 scenes). (2) An intra-object setup focused on 21 common YCB-Video objects [4] used in pose estimation benchmarks. This configuration includes 19K training images from 370 scenes and 4K test images from 74 scenes. In addition, the dataset provides metadata and tools for researchers to create custom splits for specific research tasks.
@@ IV Experiments
We conducted experiments with two primary objectives: (1) to evaluate the effectiveness of our dataset as a training resource for grasp detection, (2) to benchmark state-of-the-art perception and grasping methods on GraspClutter6D for future research. The following sections provide a comparative analysis of grasping performance across different training datasets, followed by benchmarks for segmentation, 6D object pose estimation, and 6-DoF grasp detection.
@@ IV-A Impact of GraspClutter6D on 6-DoF Grasp Detection
| Method | Train Data | Packed | Pile | ||||||
@@
| GSR | DR | GSR | DR | GSR | DR | GSR | DR | ||
| AnyGrasp | GraspNet-1B++ [9] | 77.5 | 78.9 | 58.6 | 59.3 | 68.2 | 53.0 | 71.6 | 49.5 |
| Contact- GraspNet | ACRONYM [29] | 64.6 | 60.1 | 39.2 | 28.8 | 44.8 | 21.2 | 49.3 | 18.7 |
| GraspNet-1B [18] | 74.5 | 65.2 | 62.7 | 46.6 | 67.9 | 36.3 | 71.0 | 33.4 | |
| GraspClutter6D | 84.9 | 86.1 | 77.6 | 75.4 | 77.2 | 64.7 | 77.3 | 54.0 | |
@@
Real-world Setup. We also validated the effectiveness of GraspClutter6D in real-world robotic grasping experiments. Similar to the simulation setup, we employed two configurations (Fig. 8): 1) packed scenes with 5 objects placed upright on a table without occlusion, and 2) pile scenes where objects were randomly poured into a bin, creating naturally challenging environments at varying occlusion levels with 5, 10, and 15 objects. We used 15 standardized YCB objects [37] as test items. For each grasp attempt, the highest-confidence, collision-free grasp was executed using a Robotiq 2-finger gripper mounted on a UR5 robot arm, with visual input from a RealSense D435 camera. We conducted ten rounds per configuration and method until either clearing all objects or reaching the failure threshold of 5 consecutive failures for 15-object scenes and 3 for smaller scenes.
@@
| AnyGrasp | GraspNet-1B++ [9] | 84.5 (49/58) | 54.4 (37/68) | 57.1 (72/126) | 58.4 (122/209) |
| Contact- GraspNet | GraspNet-1B [7, 18] | 77.0 (47/61) | 67.2 (39/58) | 55.9 (85/152) | 51.1 (114/223) |
| GraspClutter6D | 86.2 (50/58) | 76.9 (50/65) | 73.4 (94/128) | 68.5 (137/200) | |
@@
Real-world Results. Table IV and Fig. 9 present a comparison of grasping performance in the real world. We observed that Contact-GraspNet trained on GraspClutter6D consistently outperformed other methods in all cases. Notably, the network trained on our dataset outperformed AnyGrasp in 15 object scenes (68.5% and 58.4%), highlighting GraspClutter6D as a valuable training source for grasping in complex cluttered environments. However, achieving robust grasping in highly cluttered scenes remains a significant challenge, offering an exciting direction for future research.
IV-B Instance Segmentation Benchmarks
@@
Setup. We evaluated instance segmentation performance on unknown objects in cluttered scenes using state-of-the-art models: Mask R-CNN [1], Cascade Mask R-CNN [50], and Mask2Former [2]. All models utilized a ResNet-50 FPN [51, 52] backbone with the standard 1 training schedule and were trained and tested on GraspClutter6D using a cross-object setup, where training and test sets contain different objects. We also evaluated Grounded-SAM [53], a generalist foundation model trained on web-scale datasets, on GraspClutter6D without fine-tuning. For Grounded-SAM, we used a consistent prompt (“an object”) and filtered out detections larger than half the image size to reduce false positives. We report standard COCO metrics [54]: precision at thresholds of 0.5 (AP50), 0.75 (AP75), and average precision and recall (AP and AR) across thresholds from 0.5 to 0.95.
@@
Results. Table V presents segmentation performance on GraspClutter6D. Mask2Former achieved the highest AP of 43.5, outperforming other models. Grounded-SAM demonstrated the highest recall of AR 55.3, but its low precision (AP of 16.2%) reveals limitations in accurately delineating instance boundaries and distinguishing between individual objects in cluttered environments. These results suggest that domain-specific trained models remain more effective for segmentation tasks in complex environments, though foundation models with appropriate adaptation strategies merit further exploration for robotic applications.
IV-C 6D Object Pose Estimation Benchmarks
@@
Setup. We evaluated object pose estimation on the standard 21 YCB-Video objects [37, 4] in GraspClutter6D using two types of approaches. First, models specialized for known objects—FFB6D [5] and GDR-Net [55]—were trained on 19K images (intra-object setup) with their default configurations. Second, foundation models designed for novel object generalization—MegaPose [32] and FoundationPose [6]—were evaluated, which were trained on large-scale synthetic datasets (2M and 1.2M images, respectively).
We assessed performance using the ADD(-S) metric (point-to-point distance for non-symmetric objects, point-to-nearest-point for symmetric ones) and the ADD-S metric (point-to-nearest-point distance for all objects) [56, 57], with ADD(-S) being the stricter metric. Following established protocols [4], we computed the area under the accuracy-threshold curve up to 10 cm. Results are categorized by occlusion levels based on instance visibility : low (), medium (), and high occlusion (). For fair comparison, all methods utilized 2D detections from Mask R-CNN [1], except FFB6D which jointly performs segmentation and pose estimation.
@@
Results. Table VI presents 6D object pose estimation performance on GraspClutter6D. Foundation models substantially outperformed domain-specific approaches, with FoundationPose and MegaPose achieving 70.5 and 69.6 ADD(-S), respectively, compared to 44.5 and 36.2 for GDR-Net and FFB6D. These findings suggest that training on large-scale data is effective in improving generalizability for object pose estimation. All methods showed consistent performance degradation with increasing occlusion levels, with models experiencing substantial drops between low and high occlusion levels (e.g., MegaPose: 77.9 to 42.3 ADD(-S)). This pattern indicates that occlusion handling remains a persistent challenge in cluttered environments.
IV-D 6-DoF Grasp Detection Benchmarks
@@
Setup. We evaluated four grasping methods: Contact-GraspNet [8], GraspNet-Baseline [7], ScaleBalancedGrasp [26], and EconomicGrasp [27]. All models were trained on the standard GraspNet-1B [7, 18] with 100 train scenes and evaluated on (1) GraspNet-1B with 90 test scenes and (2) GraspClutter6D with 235 test scenes, cross-object setup. We compute the standard average precision (APμ) [18], which measures the average success rate of the top 50 predicted grasps at friction coefficient using the force-closure metric [46, 25]. We report AP as a primary metric, averaging APμ across friction coefficients from 0.2 to 1.2 at 0.2 intervals. A RealSense D435 camera was used in all datasets.
@@
Results. Table VII compares the performance of grasping methods across datasets. All methods demonstrate significant performance degradation on GraspClutter6D compared to GraspNet-1B, with EconomicGrasp exhibiting a 32.61 AP reduction. This performance gap indicates that GraspClutter6D can serve as a challenging benchmark for cluttered scene grasping. ScaleBalancedGrasp performs slightly better than EconomicGrasp on GraspClutter6D, primarily due to its auxiliary segmentation network that eliminates false positive grasps on background elements such as shelves. The consistent performance drop across all methods indicates substantial room for improvement in real-world cluttered grasping scenarios, revealing limitations in current techniques for handling complex arrangements and environments.

### v3

@@ Abstract
Robust grasping in cluttered environments remains an open challenge in robotics. While benchmark datasets have significantly advanced deep learning methods, they mainly focus on simplistic scenes with light occlusion and insufficient diversity, limiting their applicability to practical scenarios. We present GraspClutter6D, a large-scale real-world grasping dataset featuring: (1) 1,000 highly cluttered scenes with dense arrangements (14.1 objects/scene, 62.6% occlusion), (2) comprehensive coverage across 200 objects in 75 environment configurations (bins, shelves, and tables) captured using four RGB-D cameras from multiple viewpoints, and (3) rich annotations including 736K 6D object poses and 9.3B feasible robotic grasps for 52K RGB-D images. We benchmark state-of-the-art segmentation, object pose estimation, and grasp detection methods to provide key insights into challenges in cluttered environments. Additionally, we validate the dataset’s effectiveness as a training resource, demonstrating that grasping networks trained on GraspClutter6D significantly outperform those trained on existing datasets in both simulation and real-world experiments. The dataset, toolkit, and annotation tools are publicly available on our project website: https://sites.google.com/view/graspclutter6d.
I INTRODUCTION
@@
Improved grasping and perception: Models trained on our dataset demonstrate significantly better performance than those trained on existing benchmark datasets.
@@ II RELATED WORK
Robotic Grasping Datasets. State-of-the-art robotic grasping methods rely on deep learning approaches, which necessitate large-scale, diverse datasets for robust performance. Early datasets such as Cornell [35] and Jacquard [36] provided valuable resources for grasp learning. However, they were limited to planar grasping in single-object scenes. OCID-Grasp [22] extended this paradigm to cluttered scenes but offered only 75K planar grasp annotations. Synthetic datasets can be scalable alternatives: DexNet [25] with 6.7M planar grasps, ACRONYM [29] with 6-DoF grasps for 8K objects, and MetaGraspNetv2 [14] with 8,000 synthetic scenes. However, they face significant sim-to-real gaps, such as inaccurate sensor noise and contact dynamics. Among real-world 6-DoF grasp datasets, only a few public datasets exist. GraspNet-1B [7, 18], a widely adopted grasping benchmark, provides 1.1B grasp poses across 190 scenes. Though it established a foundation for 6-DoF grasping research, it remains confined to simple green tabletop environments with relatively low complexity (35.2% occlusion) and no background variation. The recent HouseCat6D dataset [19] enhances diversity with 194 objects, yet remains limited in scale and variation, offering only 10M grasps for 41 table-only scenes with low occlusion (23.5%). GraspClutter6D bridges these gaps by providing 9.3B 6-DoF grasp annotations across 1,000 highly cluttered scenes.
@@ III-A Data Acquisition
Camera Calibration. We performed extensive calibration for high-quality RGB-D data. We first conducted intrinsic calibration using a ChArUco board [40]. Then, relative poses between cameras were obtained using MC-Calib [41], followed by bundle adjustment with PnP-RANSAC [42] applied to matched points [43, 44] on key frames. Camera poses relative to the robot were determined using ArUco markers [40]. To address systematic depth errors in low-cost RGB-D cameras, we applied depth correction through least-squares fitting with linear models [12]. This reduced mean absolute depth errors from to for the RealSense D415, from to for the RealSense D435, and from to for the Azure Kinect. Our dataset provides undistorted, depth-corrected RGB-D pairs with camera poses, requiring no additional post-processing.
@@ III-B Annotation Pipelines
We first generated integrated point clouds of each scene by merging data from 13 different Zivid camera views. Annotators then aligned 3D object models to these sub-millimeter accurate point clouds, targeting a mean depth error below per object. Our tool visualized depth errors between annotated object models and scene point clouds, allowing annotators to monitor quality and refine alignments through additional adjustment and ICP [45] refinement upon saving each annotation. Independent experts verified all annotations through a double-review process. The verified poses were propagated to all camera views using camera intrinsic parameters and extrinsic poses. This process yielded annotated instances with 6D object poses and segmentation masks. The annotation tool is available on our project website.
@@ III-C Dataset Statistics
Dataset Splits. For standardized evaluation, we provide two main splits: (1) A cross-object setup that tests generalization to novel objects. This split uses 68 unseen YCB-HOPE objects [37, 38] for testing (12K images, 235 scenes), widely used robotic research benchmark objects, and 132 objects for training (32K images, 413 scenes). (2) An intra-object setup focused on 21 common YCB-Video objects [4] used in pose estimation benchmarks. This configuration includes 19K training images from 370 scenes and 4K test images from 74 scenes. The dataset also provides metadata for researchers to create custom splits for specific tasks.
@@ IV Experiments
We conducted experiments with two primary objectives: (1) to evaluate the effectiveness of our dataset as a training resource for grasp detection and perception models, (2) to benchmark state-of-the-art methods on GraspClutter6D for future research. The following sections provide a comparative analysis of grasping performance across different training datasets, followed by benchmarks for segmentation, 6D object pose estimation, and 6-DoF grasp detection.
@@ IV-A Impact of GraspClutter6D on 6-DoF Grasp Detection
| Method | Train Data | Packed | Pile | ||||||||
@@
| GSR | DR | GSR | DR | GSR | DR | GSR | DR | ||||
| AnyGrasp | GraspNet-1B++ [9] | 77.5 | 78.9 | 58.6 | 59.3 | 68.2 | 53.0 | 71.6 | 49.5 | ||
ACRONYM [29] | 64.6 | 60.1 | 39.2 | 28.8 | 44.8 | 21.2 | 49.3 | 18.7 | ||
| GraspNet-1B [18] | 74.5 | 65.2 | 62.7 | 46.6 | 67.9 | 36.3 | 71.0 | 33.4 | |||
| GraspClutter6D | 84.9 | 86.1 | 77.6 | 75.4 | 77.2 | 64.7 | 77.3 | 54.0 | |||
@@
Real-world Setup. We validated the effectiveness of GraspClutter6D in real-world robotic grasping. Similar to the simulation setup, we employed two configurations (Fig. 8): 1) packed scenes with 5 objects placed upright on a table without occlusion, and 2) pile scenes where objects were randomly poured into a bin, creating challenging environments at varying occlusion levels with 5, 10, and 15 objects. We used 15 unseen YCB objects [37] as test items, ensuring no overlap with any training sets for fair assessment. For each attempt, models first predict grasps from the RealSense D435. Then we apply collision detection [7, 18] and background filtering (5cm threshold from segmented foreground objects using [31]). We executed the highest-confidence grasp among the remaining grasps using the Robotiq gripper on a UR5. For 5-object scenes, we standardized arrangements to minimize variance across methods, while for scenes with more than 5 objects, we used identical object sequences with random placement. We conducted 20 rounds for each method and setup, setting maximum consecutive failures at 3 (10 objects) and 5 (15 objects) per round.
@@
| AnyGrasp | GraspNet-1B++ [9] | 62.7(84/134) | 66.4(81/122) | 60.7(131/216) | 59.6(252/423) |
| Contact- GraspNet | ACRONYM [29] | 32.1(34/106) | 27.8 (25/90) | 34.2 (42/123) | 25.7 (46/179) |
| GraspNet-1B [7, 18] | 77.5(93/120) | 62.5(80/128) | 58.8(124/211) | 54.9(230/419) | |
| GraspClutter6D | 93.4(99/106) | 77.2(95/123) | 74.0(174/235) | 67.9(287/423) | |
@@
Real-world Results. Table IV and Fig. 9 present a comparison of grasping performance in the real world. We observed that Contact-GraspNet trained on GraspClutter6D consistently outperformed other methods in all cases. Notably, the network trained on our dataset outperformed AnyGrasp in 15-object scenes (67.9% and 59.6%), highlighting GraspClutter6D as a valuable training source for grasping in complex cluttered environments. However, achieving robust grasping in highly cluttered scenes remains a significant challenge, offering an exciting direction for future research.
IV-B Impact of GraspClutter6D on Robotic Perception
Setup. We evaluated the effectiveness of GraspClutter6D as a training resource for perception tasks in comparison with GraspNet-1B. We used the standard YCB-Video [4] dataset as our test set (21K images) and trained Mask2Former [2] for segmentation and FFB6D [5] for 6D pose estimation using three different training configurations: (1) YCB-Video only (110K images), (2) YCB-Video + GraspNet-1B (56K additional images), and (3) YCB-Video + GraspClutter6D (14K additional images). We followed the standard evaluation protocols using COCO [50] metrics for segmentation, and ADD metrics [4] for pose estimation.
Results. Tables V and VI present segmentation and pose estimation performance on YCB-Video across different training configurations. Models trained with GraspClutter6D yield substantially greater performance improvements despite using 4 fewer additional training images (14K vs 56K). These results demonstrate that our dataset provides enhanced training benefits through increased complexity and diversity.
IV-C Instance Segmentation Benchmarks
@@
Setup. We evaluated segmentation performance on unknown objects in cluttered scenes using state-of-the-art models: Mask R-CNN [1], Cascade Mask R-CNN [51], and Mask2Former [2]. All models utilized a ResNet-50 FPN [52, 53] backbone with the standard 1 training schedule and were trained and tested on GraspClutter6D using a cross-object setup, where training and test sets contain different objects. We also evaluated Grounded-SAM [54], a generalist foundation model trained on web-scale datasets, on GraspClutter6D without fine-tuning. For Grounded-SAM, we used a consistent prompt (“an object”) and filtered out detections larger than half the image size to reduce false positives. We report average precision and recall using COCO metrics [50].
@@
Results. Table VII presents segmentation performance on GraspClutter6D. Mask2Former achieved the highest AP of 43.5, outperforming other models. Grounded-SAM demonstrated the highest recall of AR 55.3, but its low precision (AP of 16.2%) reveals limitations in accurately delineating instance boundaries and distinguishing between individual objects in cluttered environments. These results suggest that domain-specific models remain more effective for segmentation in complex environments, though foundation models with adaptation strategies merit further exploration.
IV-D 6D Object Pose Estimation Benchmarks
@@
Setup. We evaluated object pose estimation on the standard 21 YCB-Video objects [37, 4] in GraspClutter6D using two types of approaches. First, models specialized for known objects—FFB6D [5] and GDR-Net [55]—were trained on 19K images (intra-object setup) with their default configurations. Second, foundation models designed for novel object generalization—MegaPose [32] and FoundationPose [6]—were evaluated, which were trained on large-scale synthetic datasets (2M and 1.2M images, respectively). We assessed performance using the ADD(-S) and the ADD-S metric [56, 57]. Following established protocols [4], we computed the area under the accuracy-threshold curve up to 10 cm. Results are categorized by occlusion levels based on instance visibility : low (), medium (), and high occlusion (). For fair comparison, all methods utilized 2D detections from Mask R-CNN [1], except FFB6D which jointly performs segmentation and pose estimation.
@@
Results. Table VIII presents object pose estimation results on GraspClutter6D. Foundation models substantially outperformed domain-specific methods. These findings suggest that training on large-scale data is effective in improving generalizability for object pose estimation. All methods showed consistent performance degradation with increasing occlusion levels (e.g., MegaPose: 77.9 to 42.3 ADD(-S)), indicating that occlusion handling remains a persistent challenge.
IV-E 6-DoF Grasp Detection Benchmarks
@@
Setup. We evaluated four methods: Contact-GraspNet [8], GraspNet-Baseline [7], ScaleBalancedGrasp [26], and EconomicGrasp [27]. All models were trained on GraspNet-1B [7, 18] with 100 train scenes and evaluated on (1) GraspNet-1B with 90 test scenes and (2) GraspClutter6D with 235 test scenes, cross-object setup. We compute the standard average precision (APμ) [18], which measures the average success rate of the top 50 predicted grasps at friction coefficient using the force-closure metric [46, 25]. We report AP as a primary metric, averaging APμ across friction coefficients from 0.2 to 1.2 at 0.2 intervals. A RealSense D435 camera was used in all datasets. To remove background-induced ambiguity, all models used workspace-cropped inputs and were evaluated for only foreground grasps (5cm from target objects).
@@
Results. Table IX compares the performance of grasping methods across datasets. All methods demonstrate significant performance degradation on GraspClutter6D compared to GraspNet-1B, with EconomicGrasp exhibiting a 29.96 AP reduction. This performance gap indicates that GraspClutter6D can serve as a challenging grasping benchmark, highlighting substantial room for improvement in real-world clutter.